At a concession stand five hot dogs and four hamburgers cost $11.00; four hot dogs and five hamburgers cost $11.50. Find the cost of one hot dog and the cost of one hamburger.

Answer by brother Peterson Abu(1)

from the question,let X=hot dogs and Y=hamburgers
then arranging them in simultaneous equations will
give- 5X+4Y=11.......(1)
4X+5Y=11.5.....(2)
solving them with either elimination method or
substitution method will give you X=1 or $1 and Y=1.5 or $1.5.....